The length of a rectangle exceeds its width by 4 inches. find the dimensions of the rectangle if the area is 12 square inches
Answer by math_tutor2020(3816)


w = width
w+4 = length

width*length = area
w*(w+4) = 12
w^2+4w = 12
w^2+4w-12 = 0

Use the quadratic formula to solve for w.
Or you can guess-and-check to factor.
Either way you should find that:
w = -6, w = 2
Ignore the negative value. A negative width is not possible.

If w = 2, then w+4 = 2+4 = 6
The rectangle is 2 inches by 6 inches.